# A Non-perturbative General Approximation Scheme (NGAS) for Interacting Quantum Systems with Perturbation Theory for Arbitrary Strength of Interaction

### v2

@@ XIV.1 The Quartic- and Sextic Anharmonic Oscillators (AHO)
which is the Hamiltonian of an energy-shifted-harmonic-oscillator involving two free-parameters:$h_{0}$ and $\omega$, corresponding to an ‘energy-shift’ and the ‘frequency’ respectively. These parameters are determined [82] in accordance to the procedure through the steps(i)-(iii)as outlined in Section-VIA above. For the case of the QAHO, these are determined by equations(29-30) but reproduced here below for ready reference:
[…]
@@
[…]
where, the notation is : $\xi\equiv(n+\frac{1}{2});n=0,1,2,3,...$, being the spectral-level index. As derived earlier (in Section-VI-A), these parameters acquire the required functional dependence on $g$ and $\xi$:${}\omega$=$\omega(g,\xi$); ${}h_{0}=h_{0}(g,\xi)$. Similarly, the eigenvalue of the MFH, $E_{0}$ is obtained as a function of $g$ and $\xi$ and given by eqn.(32) (reproduced here, for ready reference)
[…]
(Let us note that because of the imposed constraint of CEQA ( see, Section-VI, eqn.(16)), one can also designate $H_{0}$ as the ”Mean Field Hamiltonian (MFH)”). In Section-VIIB, it has been already discussed that the leading-order(LO) result given by eqn.(32), provides [82] constitutes the dominant contribution to the energy-spectra over of the QAHO over the entire physical range of $n$ and $g$. Besides, it also reproduces the rigorous results [11] on analytic-properties of $E(g,\xi)$ in the complex-$g$ plane.
The next task is to develop the ”mean field perturbation theory (MFPT)” for the QAHO by defining the” perturbation-Hamiltonian as given generally by eqn.(21). For the case of the QAHO. this yields:
@@
For the case of the QAHO, we note the following results for $E_{p}(g,\xi)$ at $g=1$ and $\xi=\frac{1}{2}$ (i.e. for the ground-state), for some low-lying values of $p$: $E_{1}=0$, $E_{2}=-\frac{3}{256}$,$E_{3}=\frac{27}{4096}$, $E_{4}=-\frac{2373}{262144}$, $E_{5}=\frac{65457}{4194304}....$ etc. (Note that consistent with eqn.(20), we recover the result :$E_{1}=0$ in the limit: $\eta~{}\rightarrow~{}1~{}$, which is of special significance- see later. )
@@
Some sample-values of $E_{p}$ for the ground state i.e.$\xi=\frac{1}{2}$ and for $\omega=2$ are the following: $E_{1}=0$, $E_{2}=-\frac{49}{960}$, $E_{3}=\frac{671}{4608}$, $E_{4}=-\frac{53621891}{55296000}$, $E_{5}=\frac{2610955409}{265420800}....$ etc. As expected [100,102], the resultant asymptotic series exhibits more severe divergence at large orders as compared to that of the QAHO.
@@ XIV.2 The Quartic Double Well Oscillator (QDWO)
The QDWO constitutes perhaps the simplest case displaying degeneracy of the ground-state. As such, it has played a very important role in the elucidation [82,95,98] of the properties of standard formulation of perturbation theory (SFPT) . In particular,as has been mentioned earlier, successive terms in the perturbation series (PS) of SFPT for the energy eigenvalue are found to be non-alternating in sign at large orders, which frustrates Borel-summability [82,95,98].Several modern developments such as the theory of resurgence [95-99], trans-series [97,98] etc are basically motivated to surmount such typical problems of Borel-nonsummability arising in SFPT when dealing with systems having degenerate ground states.In view of the above scenario,we consider it specially relevant to investigate the QDWO-problem in the frame-work of MFPT.
The Hamiltonian in this case, is given by (see, Section-VIIC):
[…]
@@
[…]
This choice is motivated to account for the spontaneous symmetry breaking (SSB) through a non-zero vacuum-expectation-value (VEV) of the field operator $\hat{x}$, denoted here by $\sigma$. The parameters $\omega$, $\sigma$ and $h_{0}$ are determined in analogous manner as in the case of the QAHO. The LO-results [82] (See, Section-VIIC) imply ”quantum-phase transition’(QPT) characterized by a critical-coupling:$g_{c}(\xi)$ given by the expression:
@@
[…]
and  $h_{0}=E_{0}-\omega\xi$
It has been demonstrated in Section-VII and ref. [82] that LO-results for $E_{0}$ as given above, provide accurate estimate of the energy-spectrum to within a few-percent of the ‘exact’-result for arbitrary value of $g$ and $\xi$, thus constituting the major contribution to it. Further improvement over the LO-result can be attempted by the application of MFPT following the procedure analogous to that for the QAHO. For this purpose, the ’perturbation-Hamiltonian’ and the ’associated- Hamiltonian’ are now respectively given by:
@@
As stated earlier (Section-VIIC), a quantum- phase-transition(QPT) to this phase occurs when $g<g_{c}(\xi)$ manifesting in degenerate global-minima of the effective potential located at $x=\pm\sigma$, where ${}~{}~{}\sigma^{2}=(1-{12g{\xi}}/{\omega})/4g.$ In this case, $\omega$ satisfies[82] the equation,
${}~{}~{}~{}\omega^{3}-2\omega+6g(5\xi-\frac{1}{4\xi})=0.$ (225)
@@
${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}E_{0}={(-1/16g)}+{(\xi/4)}{(3\omega+(2/\omega))}~{}$
and the energy-shift is given by: ${}h_{0}=E_{0}-\omega\xi$. Again, the LO-result for $E_{0}$ as given above, provides [82] accurate estimate of the energy-spectrum to within a few-percent of the ‘exact’-numerical result for arbitrary values of $g$ and $\xi$. Further improvement in accuracy can be achieved by computing perturbation-corrections to the LO-result via MFPT as described below: The perturbation-Hamiltonian and the auxiliary Hamiltonian (AH) are respectively given by,
@@
Due to the presence of odd-powers of $x$ in the expression of $\bar{H}$ as given above, it is appropriate to choose:$f(x)=x^{j},j=0,1,2,3,..$ for the application of HVT to $\bar{H}$. This leads to the following recursion relation [110,111]:
@@
where, $X(j)\equiv\langle\bar{\psi}|x^{j}|\bar{\psi}\rangle$,${}X(j,i)$ are defined by the relation: $X(j)=\sum\limits_{i=0}^{\infty}{\eta^{i}}X(j,i)$
@@
[…]
@@ XV Computation of the Total Perturbation Correction ( TPC )
We compute the TPC by two main methods: (a) the method of ‘optimal truncation’ (MOT) of the original asymptotic perturbation series and (b) by ’Borel-summation’. These are sequentially described in the following sections. (For convenience and illustration, we confine here to the computation for the ground state.)
@@ XV.1 Method of Optimal Truncation (MOT)
In this Table-I, we have listed computed values for the ’total perturbation correction (TPC) ’ by the method of ’optimal truncation(MOT)’ of perturbation series(APS) in MFPT, for the ground-state energy at different values of the coupling-strength $g$ (col.(1)). Results are shown (col.(4)) for the ground-state energy of QAHO, SAHO and QDWO and compared with accurate numerical results of earlier calculations denoted as `$Exact^{\prime}$ (col.(5)), from the given reference. The LO-results in MFPT are given in col.(2). In col.3, $N_{0}(g)$ denotes the ”truncation-point” of the respective APS at the‘term of least magnitude’(TLM). Accuracy as percentage-errors are depicted in the last-column relative to the‘exact’-results. For comparison with ‘standard’ results, we have considered earlier computations in [112],[113] and [114] for the case of QAHO, SAHO and QDWO respectively. The relative-error incurred is seen to be within a few-percent uniformly over the considered range of $g$.
@@
$QAHO$ $ref.(112)$ $g$ $E_{0}$ $N_{0}(g)$ $E_{MOT}$ $Exact$ $Error{(\%)}$ 0.1 0.5603 6 0.5593 0.5591 0.03 1.0 0.8125 3 0.8074 0.8038 0.44 10.0 1.5312 3 1.5204 1.5050 1.02 100.0 3.1924 3 3.1701 3.1314 1.23 $SAHO$ $ref.(113)$ 0.1 0.5964 2 0.5787 0.5869 1.40 1.0 0.8378 2 0.7694 0.8050 4.42 50.0 1.9735 2 1.7241 1.8585 7.23 200.0 2.7606 2 2.3986 2.5942 7.54 $QDWO$ $ref.(114)$ 0.1 0.5496 2 0.4107 0.4709 12.78 1.0 0.5989 3 0.5998 0.5773 3.91 10.0 1.4097 3 1.4007 1.3778 1.66 100.0 3.1338 3 3.1122 3.0701 1.37
@@ XV.2 Method of Borel Summation (MBS)
[…]
where, we have defined,${}u=t^{\alpha},\gamma=\frac{1}{\alpha}$ and $B(u)$ denotes the ‘Borel-Series(BS)’, given by :
@@
Note that, although by virtue of eqn.(232), $B(u)$ has a finite ‘radius of convergence’ $r_{c}$ , the ‘Borel-Laplace-integral’(BLI), eqn.(233) does not exist yet since the range of integration extends beyond $r_{c}$.
The way out of this difficulty is to carry-out the analytic continuation/ extrapolation of the BS to the full-path of integration of the BLI, i.e. to the whole $+ve$-real axis in the u-plane. Denoting such analytic-continuation by $\tilde{B}(u)$ and substituting this for $B(u)$ in eqn.(234), the BLI can now be made to exist and hence can be used to‘define’ the TPC as:
[…]
@@
We obtain the value:${}p(g)=~{}-0.5$, being independent of $g$ as expected and the values for $r_{c}(g)$, as shown in Table-II. Equipped with these information as inputs, consider [118] the conformal mapping , $z$ which maps the cut-u plan,$|arg(u)|<\pi$ into the interior of the unit-disk in the $z$-plane while preserving the origin:
@@
The other important property of the conformal-transformation, eqn.(239) is its ‘optimality’ as established by the ‘Ciulli-Fischer Theorem’ [118]. According to this theorem, the error of truncation becomes minimum when an arbitrary function $f(z)$ expanded as an infinite- power-series in $z$, is approximated by truncation at a finite number of terms, thus ensuring the optimal rate of convergence.
Hence, consider the partial sum of the BS, eqn.(235):  $B_{N}(u)\equiv\sum\limits_{k=1}^{N}b_{k}u^{k}$.
Using the inverse-transformation, eqn.(239b),$B_{N}(u)$ can be consistently re-expanded [119,120] as follows, noting that $u^{N}\sim~{}z^{N}+O(z^{N+1})$:
@@
where, we have denoted the analytic-continuation of the partial sum of the BS by:${}B_{N}(u)\rightarrow~{}\bar{B}_{N}(z)$.
Using the above properties/results, the TPC can be evaluated by the change of the integration-variable to $z$ (see, eq(239b)).Then, by taking care of possible numerical-divergence at the upper limit, ${}z=1$ and further, considering the limit of the sequence of partial-sums, the TPC is evaluated as follows ( we suppress the $g$-dependence of quantities for notational-clarity):
@@
 with $\mbox{~{}with~{}}f_{N}(z)=(\gamma\rho)~{}\dfrac{(1+z)}{(1-z)^{3}}\left(\dfrac{\rho~{}z}{(1-z)^{2}}\right)^{\gamma-1}$
where,${}\rho=4r_{c}$ and ${}\bar{B}_{N}(z)$ is defined by eqn.(240a). Note that term-by-term integration of the partial-sum,${}\bar{B}_{N}(z)$ can now be carried out in eqn.(241a), in contrast to the BS in eqn.(234) because of the incorporation of analyticity through the conformal-map $z$.
@@
The results are presented in Table-II. There we have shown ( col.(4)) the Borel-sum for the total- perturbation-correction(TPC) computed using eqn.(241a-b) and conformal-mapping, eqn.(239a) , for the ground state energy of the QAHO, SAHO and the QDWO as a function of $g$ given in col.(1). For each case, the input-value of $\alpha$, (see eqn.(232)) is indicated. Entries in col.(2) are the input-value for the ‘radius of convergence’; $N_{c}$ in col.(3) denotes the number of terms retained in the Borel-series for convergence up to the specified accuracy; $E_{0}$ in col.(5) denotes the LO-contribution; $E_{tot}$ in col.(6) represents the ‘total corrected energy’; entries in col(7) labeled $Exact$, display the ‘standard’-numerical results from the indicated reference; and the last column specifies the %${\%}error$ with respect to the $Exact$-result. We have fixed the value:  $\epsilon=0.001$ ,( see eqn.(241a)) in the computation of TPC in all-cases.
It may be noted that the convergence of the partial-sum is quite fast- an accuracy of $\sim(99.99\%)$ is uniformly achieved over the full explored range of $g$ and for all the cases of anharmonic-interaction since the cut-off $N_{c}$ on the number of terms retained in the partial-sum is as low as $\sim\left(10-20\right)$ (Table-II). Thus, the ‘$optimality$’ and the ‘$accuracy$’ of the ’method of conformal mapping’ are firmly established in MFPT.

### v3

@@ XIV.1 The Quartic- and Sextic Anharmonic Oscillators (AHO)
$\omega^{3}-\omega-6g(\xi+\frac{1}{4\xi})=0,~~~~~~\mbox{(29)}$
@@
$h_{0}=\left(\frac{\xi}{4}\right)\left(\frac{1}{\omega}-\omega\right)~~~~~~~~~~\mbox{(30)}$
where, the notation is : $\xi\equiv(n+\frac{1}{2});n=0,1,2,3,$..., being the spectral-level index. As derived earlier (in Section-VI-A), these parameters acquire the required functional dependence on $g$ and $\xi$$\omega$$\omega(g,\xi$); $h_{0}=h_{0}(g,\xi)$. Similarly, the eigenvalue of the MFH, $E_{0}$ is obtained as a function of $g$ and $\xi$ and given by eqn.(32) (reproduced here, for ready reference)
$E_{0}(g,\xi)=\left(\frac{\xi}{4}\right)\left(3\omega+\frac{1}{\omega}\right).~~~~\mbox{(32)}$
(Let us note that because of the imposed constraint of CEQA ( see, Section-VI, eqn.(16)), one can also designate $H_{0}$ as the ”Mean Field Hamiltonian (MFH)”). In Section-VIIB, it has been already discussed that the leading-order(LO) result given by eqn.(32), provides [82] the dominant contribution to the energy-spectra over of the QAHO over the entire physical range of $n$ and $g$. Additionally, it reproduces the rigorous results [11] on the analytic properties of $E(g,\xi)$ in the complex-$g$ plane.
The next task is to develop the ”mean field perturbation theory (MFPT)” for the QAHO by defining the ”perturbation Hamiltonian” as generally given by eqn. (21). For the case of the QAHO, this yields:
@@
For the case of the QAHO, we note the following results for $E_{p}(g,\xi)$ at $g=1$ and $\xi=\frac{1}{2}$ (i.e. for the ground-state), for some low-lying values of $p$: $E_{1}=0$, $E_{2}=-\frac{3}{256}$$E_{3}=\frac{27}{4096}$, $E_{4}=-\frac{2373}{262144}$, $E_{5}=\frac{65457}{4194304}.$... etc. (Note that consistent with eqn.(20), we recover the result $E_{1}=0$ in the limit: $\eta~\rightarrow~1$~, which is of special significance- see later. )
@@
[…]
@@ XIV.2 The Quartic Double Well Oscillator (QDWO)
The QDWO represents perhaps the most straightforward example exhibiting ground-state degeneracy. Consequently, it has been instrumental in elucidating [82,95,98] the characteristics of the standard formulation of perturbation theory (SFPT). Notably, as previously mentioned, the successive terms in the perturbation series (PS) of SFPT for the energy eigenvalue are non-alternating in sign at higher orders, which impedes Borel-summability [82,95,98]. Several contemporary advancements, such as the theory of resurgence [95-99] and trans-series [97,98], are fundamentally motivated to overcome such typical issues of Borel-nonsummability encountered in SFPT when addressing systems with degenerate ground states. In light of this context, we find it particularly pertinent to examine the QDWO problem within the framework of MFPT.
The Hamiltonian in this instance is expressed as (see, Section-VIIC):
$H~=~\frac{1}{2}p^{2}-\frac{1}{2}x^{2}+gx^{4},~~~~~~~~~~~~~~\mbox{(36)}$
@@
$H_{0}~=~\frac{1}{2}p^{2}+\frac{1}{2}\omega^{2}(x-\sigma)^{2}+h_{0}.~~~~~~\mbox{(37)}$
This choice is motivated to account for the spontaneous symmetry breaking (SSB) through a non-zero vacuum-expectation-value (VEV) of the field operator $\hat{x}$, denoted here by $\sigma$. The parameters $\omega$, $\sigma$, and $h_{0}$ are determined analogously to the case of the QAHO. The LO-results [82] (See, Section-VIIC) imply a ”quantum-phase transition” (QPT) characterized by a critical coupling: $g_{c}(\xi)$ given by the expression:
@@
$E_{0}=(\xi/4)(3\omega-(1/\omega));~~~~~~~~~~~~~~~~~~~~~~~~~\mbox{(42)};$
and  $h_{0}=E_{0}-\omega\xi$ .
It has been demonstrated in Section-VII and ref. [82] that LO-results for $E_{0}$ as given above, provide an accurate estimate of the energy spectrum to within a few percent of the ’exact’ result for arbitrary values of $g$ and $\xi$, thus constituting the major contribution to it. Further improvement over the LO-result can be attempted by the application of MFPT following the procedure analogous to that for the QAHO. For this purpose, the ’perturbation-Hamiltonian’ and the ’associated-Hamiltonian’ are now respectively given by:
@@
As stated earlier (Section-VIIC), a quantum- phase-transition(QPT) to this phase occurs when $g<g_{c}(\xi)$ manifesting in degenerate global-minima of the effective potential located at $x=\pm\sigma$, where $\sigma^{2}=(1-{12g{\xi}}/{\omega})/4g.$ In this case, $\omega$ satisfies[82] the equation,
$\omega^{3}-2\omega+6g(5\xi-\frac{1}{4\xi})=0.$ (225)
@@
$E_{0}={(-1/16g)}+{(\xi/4)}{(3\omega+(2/\omega))}$~
and the energy-shift is given by: $h_{0}=E_{0}-\omega\xi$. Again, the LO-result for $E_{0}$ as given above, provides [82] accurate estimate of the energy-spectrum to within a few-percent of the ‘exact’-numerical result for arbitrary values of $g$ and $\xi$.
Further improvement in accuracy can be achieved by computing perturbation-corrections to the LO-result via MFPT as described below: The perturbation-Hamiltonian and the auxiliary Hamiltonian (AH) are respectively given by,
@@
[…]
@@
where, $X(j)\equiv\langle\bar{\psi}|x^{j}|\bar{\psi}\rangle$,X(j,i)~X(j,i) are defined by the relation: $X(j)=\sum\limits_{i=0}^{\infty}{\eta^{i}}X(j,i)$
@@
Having thus achieved the objective of analytic computation of the perturbation corrections $E_{p}$ in all the specified cases of the anharmonic-interaction, we consider the evaluation of the total $\it`sum^{\prime}$} of these corrections in the next section. In that context, the following common features that emerge from the computations in all the above cases may be noted:    $E_{1}=0$, and the perturbation series have terms that alternate in sign. These two features in MFPT are of special significance in obtaining the “sum” of the perturbation series as described below.
@@ XV Computation of the Total Perturbation Correction ( TPC )
We compute the TPC by two main methods: (a) the method of ‘optimal truncation’ (MOT) of the original asymptotic perturbation series and (b) by ’Borel-summation’. These are sequentially described in the following sections. (For convenience and definiteness , we confine here to the computation for the ground state.)
@@ XV.1 Method of Optimal Truncation (MOT)
In this Table-I, we have listed computed values for the ’total perturbation correction (TPC) ’ by the method of ’optimal truncation(MOT)’ of perturbation series(APS) in MFPT, for the ground-state energy at different values of the coupling-strength $g$ (col.(1)). Results are shown (col.(4)) for the ground-state energy of QAHO, SAHO and QDWO and compared with accurate numerical results of earlier calculations denoted as $Exact^{\prime}$ (col.(5)), from the given reference. The LO-results in MFPT are given in col.(2). In col.3, $N_{0}(g)$ denotes the ”truncation-point” of the respective APS at the‘term of least magnitude’(TLM). Accuracy as percentage-errors are depicted in the last-column relative to the‘exact’-results. For comparison with ‘standard’ results, we have considered earlier computations in [112],[113] and [114] for the case of QAHO, SAHO and QDWO respectively. The relative-error incurred is seen to be within a few-percent uniformly over the considered range of $g$.
@@
QAHOQAHO $ref.(112)$ $g$ $E_{0}$ $N_{0}(g)$ $E_{MOT}$ ExactExact $Error{(\%)}$ 0.1 0.5603 6 0.5593 0.5591 0.03 1.0 0.8125 3 0.8074 0.8038 0.44 10.0 1.5312 3 1.5204 1.5050 1.02 100.0 3.1924 3 3.1701 3.1314 1.23 SAHOSAHO $ref.(113)$ 0.1 0.5964 2 0.5787 0.5869 1.40 1.0 0.8378 2 0.7694 0.8050 4.42 50.0 1.9735 2 1.7241 1.8585 7.23 200.0 2.7606 2 2.3986 2.5942 7.54 QDWOQDWO $ref.(114)$ 0.1 0.5496 2 0.4107 0.4709 12.78 1.0 0.5989 3 0.5998 0.5773 3.91 10.0 1.4097 3 1.4007 1.3778 1.66 100.0 3.1338 3 3.1122 3.0701 1.37
@@ XV.2 Method of Borel Summation (MBS)
$\Gamma(z)=~\int_{0}^{\infty}dt~exp(-t)t^{z-1},$ (236)
$\Delta{E(g)}\equiv\sum\limits_{k=1}^{\infty}E_{k}~=\sum\limits_{k=1}^{\infty}b_{k}\int_{0}^{\infty}dt~exp(-t)t^{\alpha_{k}}$
$=\gamma\int_{0}^{\infty}du~exp(-u^{\gamma})u^{\gamma-1}B(u),$ (237)
@@
It is important to note that, although by virtue of eqn.(232), $B(u)$ has a finite radius of convergence $r_{c}$, the Borel-Laplace-integral (BLI), eqn.(233), does not exist yet since the range of integration extends beyond $r_{c}$. The resolution to this issue is to perform the analytic continuation/extrapolation of the BS to the full path of integration of the BLI, i.e., to the entire positive real axis in the u-plane. Denoting such analytic continuation by $\tilde{B}(u)$ and substituting this for $B(u)$ in eqn.(234), the BLI can now be made to exist and hence can be used to define the TPC as:
$\Delta{E(g)}\equiv\gamma\int_{0}^{\infty}du~exp(-u^{\gamma})u^{\gamma-1}\tilde{B}(u),$ (239)
@@
[…]
@@
The conformal transformation, as described in equation (239), possesses a significant property known as its ’optimality,’ as demonstrated by the Ciulli-Fischer Theorem [118]. This theorem asserts that the truncation error is minimized when an arbitrary function $f(z)$, expressed as an infinite power series in $z$, is approximated by truncating the series to a finite number of terms, thereby ensuring an optimal rate of convergence. Consider the partial sum of the BS, as given in equation (235):
$B_{N}(u)\equiv\sum\limits_{k=1}^{N}b_{k}u^{k}$.
[…]
@@
$\mbox{where,}~~B_{k}=\sum\limits_{n=1}^{k}b_{n}\dfrac{\left(n+k-1\right)!}{\left(k-n\right)!\left(2n-1\right)!},$ (243b)
@@
Using the above properties/results, the TPC can be evaluated by the change of the integration-variable to $z$ (see, eq(239b)).Then, by taking care of possible numerical-divergence at the upper limit, z=1~z=1 and further, considering the limit of the sequence of partial-sums, the TPC is evaluated as follows ( we suppress the $g$-dependence of quantities for notational-clarity):
$\Delta{E}=\lim_{N\gg 1,~\epsilon\rightarrow 0}\left(\Delta{E}\right)_{N}\equiv\int_{0}^{1-\epsilon}dzf_{N}(z),$ (244a)
 $\mbox{~with~}f_{N}(z)=(\gamma\rho)~\dfrac{(1+z)}{(1-z)^{3}}\left(\dfrac{\rho~z}{(1-z)^{2}}\right)^{\gamma-1}$
@@
The results are presented in Table-II. There we have shown ( col.(4)) the Borel-sum for the total- perturbation-correction(TPC) computed using eqn.(241a-b) and conformal-mapping, eqn.(239a) , for the ground state energy of the QAHO, SAHO and the QDWO as a function of $g$ given in col.(1). For each case, the input-value of $\alpha$, (see eqn.(232)) is indicated. Entries in col.(2) are the input-value for the ‘radius of convergence’; $N_{c}$ in col.(3) denotes the number of terms retained in the Borel-series for convergence up to the specified accuracy; $E_{0}$ in col.(5) denotes the LO-contribution; $E_{tot}$ in col.(6) represents the ‘total corrected energy’; entries in col(7) labeled $Exact$, display the ‘standard’-numerical results from the indicated reference; and the last column specifies the ${\%}error$ with respect to the $Exact$-result. We have fixed the value:  $\epsilon=0.001$ ,( see eqn.(241a)) in the computation of TPC in all-cases.
It is noteworthy that the convergence of the partial-sum is quite fast- an accuracy of $\sim(99.99\%)$ is uniformly achieved over the full explored range of $g$ and for all the cases of anharmonic-interaction since the cut-off $N_{c}$ on the number of terms retained in the partial-sum is as low as $\sim\left(10-20\right)$ (Table-II). Thus, the ‘optimalityoptimality’ and the ‘accuracyaccuracy’ of the ’method of conformal mapping’ are firmly established in MFPT.